The Integer-Sequence for the Number of Monomer-Dimer Tilings of an n by n Chessboard Starts With: 1, 7, 131, 10012, 2810694, 2989126727, 11945257052321, 179788343101980135, 10185111919160666118608, 2172138783673094193937750015, 1743829823240164494694386437970640, 5270137993816086266962874395450234534887, 59956919824257750508655631107474672284499736089, Thanks to Zeilberger’s Maple Package DOMINO ^{1}

Shalosh B. EKHAD^{2}

1 http://www.math.temple.edu/~zeilberg/tokhniot/DOMINO, [Type: MonoDimerSequence(13);] 2 Department of Mathematics, Temple University, Philadelphia, PA 19122. ekhad@math.temple.edu ;

http://www.math.temple.edu/˜ ekhad . Oct. 14, 1997. I wish to thank Steve Finch for raising the question.